Find the value of the equation 2x + 3y = -16
5x- 10x = 30 in substitutional method

Found 3 solutions by Theo, greenestamps, ikleyn:
Answer by Theo(13342)

i think you meant 5x - 10y rather than 5x - 10x.
otherwise the second equation doesn't make much sense.
i'll assume 5x - 10y when i solve.
if, for some reason, you really did mean 5x - 10x, let me know.

your equations are, based on my assumption:

2x + 3y = -16
5x - 10y = 30

solve for x in the second equation to get:
5x - 10y = 30
add 10y to both sides of the equation to get:
5x = 10y + 30
solve for x to get:
x = 2y + 6

replace x in the first equation with that to get:
2x + 3y = -16 becomes:
2 * (2y + 6) + 3y = -16 which becomes:
4y + 12 + 3y = -16
combine like terms to get:
7y + 12 = -16
subtract 12 from both sides of the equation to get:
7y = -28
solve for y to get:
y = -4

replace y with -4 in the first equation to get:
2x + 3y = -16 becomes:2x - 12 = -16
add 12 to both sides of the equation to get:
2x = -4
solve for x to get:
x = -2

you have y = -4 and x = -2

first equation of 2x + 3y = -16 becomes:
2 * -2 + 3 * -4 = -16 which becomes:
-4 - 12 = -16 which becomes:
-16 = -16 which is true.

second equation of 5x - 10y = 30 becomes:
5 * -2 - 10 * -4 = 30 which becomes:
-10 + 40 = 30 which becomes:
30 = 30 which is true.

your solution is x = -2 and y = -4.

if i assume that you really meant 5x - 10x = 30, then:
your equations are:
2x + 3y = -216
5x - 10x = 30

in the second equation, combine like terms to get -5x = 30
solve for x to get:
x = -6

in the first equation of 2x + 3y = -16, replace x with -6 to get:
2 * -6 + 3y = -16
simplify to get:
-12 + 3y = -16
add 12 to both sides of the equation to get:
3y = -4
solve for y to get:
y = -4/3

in the first equation of 2x + 3y = -16, replace x with -6 and y with -4/3 to get:
2x + 3y = -16 becomes:
2 * -6 + 3 * -4/3 = -16 which becomes:
-12 -4 = -16 which becomes:
-16 = -16 which is true.

your solution is x = -6 and y = -4/3


Answer by greenestamps(13203)



2x + 3y = -16
5x - 10y = 30

Simplify the second equation and solve for x:

x - 2y = 6
x = 2y + 6

Substitute for x in the first equation and solve for y:

2(2y + 6) + 3y = -16
4y + 12 + 3y = -16
7y = -28
y = -4

Substitute in either original equation and solve for x:

2x + 3(-4) = -16
2x - 12 = -16
2x = -4
x = -2

ANSWERS: x = -2; y = -4

You should of course check your answers in the original equations; I won't take the time to so that here.


Answer by ikleyn(52818)

.

How this problem is worded, it shows/demonstrates that the "problem's composer"
is unfamiliar with standard basic terminology in the subject (in Elementary Math).

It looks like this "problems composer" never opened and never read Math textbooks in English.